OpenStudy (goformit100):

Let N be the sum of the digits of a natural number A, let B = A+N, Let A’ be the sum of the digits of the number B, and let C = B+A’. Find A if the digits of C are those of A in reverse order.

OpenStudy (goformit100):

@amistre64

OpenStudy (amistre64):

i got nothing on this one ....

OpenStudy (goformit100):

ok sir then Finally I Must close this question :'(

OpenStudy (shubhamsrg):

B-A =N C-B = A' C-A =N + A' we know C-A is a multiple of 9 logically speaking, i'd say A is a 2 digit number as i think it is predictable that A can not be of more than 2 digits . now N + A' should be a factor of 9 assuming A= 10a+b also assuming a+2b<10 for now now B = 11a + 2b = 10a + (a+2b) C = 13a + 4b we are given 4b + 13a = 10b+a b=2a a=1 and b=2 fit in i.e A=12 it was purely accident that i got the ans :|

OpenStudy (goformit100):

Thank you sir